OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 4y = 8 slope = -one half; y-intercept = (0, 2) slope = -2; y-intercept = (0, 4) slope = -2; y-intercept = (0, 8) slope = -one half; y-intercept = (0, 8)

OpenStudy (anonymous):

Equation of line is y=mx+c where c is y intercept and m is gradient of line reduce 2x + 4y = 8 in to the form y=mx +c by dividing through with 4 and rearranging the equation to get c your y intercept

OpenStudy (itsbribro):

slope= -one half; y-intercept=(0,2) i think thats what it is!!

OpenStudy (anonymous):

so its #1????

OpenStudy (itsbribro):

yes